OpenStudy (anonymous):

Use the ideal gas law (pV = nRT) derive the relationship Vi/Ti = Vf/Tf and then Vi/Vf = Ti/Tf. Hint: hold the amount of gas, n, and the pressure, p, constant. Write two versions of the equation, an initial version and a final version, with the appropriate subscripts. In each equation get all the constants on one side and use that to connect the two equations).

OpenStudy (anonymous):

pV=nRT V/T=nR/p Vi/Ti=nR/p Vf/Tf=nR/p since n and p constant so Vi/Ti=Vf/Tf Vi/Vf=Ti/Tf

OpenStudy (anonymous):

thanks